OpenStudy (anonymous):

2x/3= -4x/5-1/2 Solve equation for x! plz!

OpenStudy (anonymous):

\[\frac{2}{3}x=-\frac{4}{5}x-\frac{1}{2}\]?

OpenStudy (anonymous):

clear the annoying fractions by multiplying both sides by 30 and get \[20x=-24x-15\]

OpenStudy (anonymous):

ohhh...

OpenStudy (anonymous):

now it should be easy right?

OpenStudy (anonymous):

yep thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):